Company XYZ manufacturers Large scale units.It has been shown that the marginal (or variable)cost, which is the gradient of the total cost curve is 92-2x where x is the number of units of outputs per annum.The fixed costs are 800000 per annum.It has also been shown that the marginal revenue which is the gradient of the total revenue is 112-2x.Establish the total cost curve,, Establish the total revenue curve,, Establish the break even situation for your company,,, Determine the number of units of output that would maximize total revenue and maximize the total cost


1
Expert's answer
2021-04-23T12:07:58-0400
"VC=92-2x"

"FC=800000"

"TC=VC+FC=800092-2x"

"TR=112+2x"

"112\n+2x=800012-2x"

"x=199975"